# Determinant of the crossing matrix of a braid

### v2

@@ Abstract
In this paper, we define a braid invariant, the purified determinant $P(b)$ of a braid $b$, considering the determinant of the crossing matrix, and show that $P(b_{1}b_{2})=P(b_{2}b_{1})$ for any pair of $n$-braids $b_{1}$ and $b_{2}$.
@@ 1 Introduction
In this paper, we introduce a new braid invariant, the purified determinant $P(b)$, of a braid $b\in\mathfrak{B}_{n}$ considering the determinant of the crossing matrix of a pure braid derived from $b$, and show that the value of $P$ is unchanged by a conjugation.
@@ 2 Crossing matrix
In this section, we review the crossing matrix and its properties. For a matrix $M$, we denote the $(i,j)$-component by $M(i,j)$. For an $n$-braid diagram $B$, the crossing matrix $C(B)=M$ was defined by Burillo, Gutierrez, Krstić and Nitecki in [5] (see also [6]) to be the $n\times n$ matrix such that $M(i,i)=0$ and $M(i,j)$ is the number of positive crossings minus negative crossings between the $i^{th}$ and $j^{th}$ strands where the $i^{th}$ strand is over the $j^{th}$ strand. (See Figure 2.)
@@
For a braid $b\in\mathfrak{B}_{n}$ or a braid diagram $B\in\mathscr{B}_{n}$, when a strand has the initial point at the $i^{th}$ position from the left and has the end point at the $j^{th}$ position from the left, we denote $\rho(i)=j$ and $\rho(1,2,\dots,n)=(j_{1},j_{2},\dots,j_{n})$ for $\rho(k)=j_{k}$. We call $\rho$ the braid permutation of $b$ or $B$. For example, the braid diagram $B$ in Figure 2 has the braid permutation $\rho$ as $\rho(1,2,3,4,5)=(2,1,4,5,3)$. We note that a braid and its diagram have the same braid permutation. By regarding $\rho$ as an element of the symmetric group, the identity $\rho^{0}=id$, the inverse $\rho^{-1}$, the exponentiation $\rho^{k}$ for $k\in\mathbb{N}$ are also defined. For a permutation $\rho$, the order, $|\rho|$, of $\rho$ is defined to be the number $k\in\mathbb{N}$ such that $\rho^{k}=id$ and $\rho^{l}\neq id$ for any $l<k$. It is well known in the study of braid group that two conjugate braids have the same order of braid permutation. We say that a braid or a braid diagram is pure when the permutation is identity. Crossing matrix of a pure braid and a positive pure braid is studied in [5], [6], [9], [10], [11].
For square matrices, we define the following operation that reorders the row and column according to a permutation $\rho$.
@@
As mentioned in [10], $M$ and $\rho(M)$ are similar, that is, there exists an invertible matrix $Q$ such that $\rho(M)=Q^{-1}MQ$, and therefore $\det(M)=\det(\rho(M))$. For two $n\times n$ matrices $M$ and $N$, we have $\rho(M+N)=\rho(M)+\rho(N)$ by definition. For the crossing matrix of a braid product, the following proposition was shown in [5].
@@ 3 Similarity of crossing matrices
Proof of Lemma 1. Let $M_{1}=C(B)$ and $M_{2}=C(B^{\prime})$, where $B^{\prime}=\sigma_{i}^{\varepsilon}B\sigma_{i}^{-\varepsilon}$. For $B^{\prime}$, the $i^{th}$ (respectively, $(i+1)^{th}$) strand corresponds to the $(i+1)^{th}$ (respectively, $i^{th}$) strand of $B$ because there is a crossing $\sigma_{i}^{\varepsilon}$ between the $i^{th}$ and $(i+1)^{th}$ strands as the first crossing of $B^{\prime}$. Also, the first crossing $\sigma_{i}^{\varepsilon}$ and the last crossing $\sigma_{i}^{-\varepsilon}$ of $B^{\prime}$ are both between the $i^{th}$ and $(i+1)^{th}$ strands because $B^{\prime}$ is pure, where the $(i+1)^{th}$ (respectively, $i^{th}$) strand is over when $\varepsilon=+1$ (respectively, $\varepsilon=-1$), and one of the two crossings is positive and the other is negative. (See Figure 3.) Hence, we have $M_{2}(i+1,i)=M_{1}(i,i+1)$ and $M_{2}(i,i+1)=M_{1}(i+1,i)$. Therefore, we have $M_{2}=\tau_{i,i+1}(M_{1})$, where $\tau_{i,i+1}$ is a permutation that swaps the $i^{th}$ and $(i+1)^{th}$ elements. This implies that $M_{1}$ and $M_{2}$ are similar to each other. ∎
@@
For non-pure braid diagrams, we have the following lemma.
@@
Let $B\in\mathscr{B}_{n}$, $i\in\{1,2,\dots,n-1\}$ and $\varepsilon\in\{+1,-1\}$. Let $r$ be the order of the permutation of the braid diagram $B$. Then, $C(B^{r})$ and $C((\sigma_{i}^{\varepsilon}B\sigma_{i}^{-\varepsilon})^{r})$ are similar to each other.
@@
Since $B^{r}$ is pure and $(\sigma_{i}^{\varepsilon}B\sigma_{i}^{-\varepsilon})^{r}=(\sigma_{i}^{\varepsilon}B\sigma_{i}^{-\varepsilon})(\sigma_{i}^{\varepsilon}B\sigma_{i}^{-\varepsilon})\dots(\sigma_{i}^{\varepsilon}B\sigma_{i}^{-\varepsilon})$ has the same crossing matrix to $\sigma_{i}^{\varepsilon}B^{r}\sigma_{i}^{-\varepsilon}$, they have similar crossing matrices by Lemma 1. ∎
Since $C(B^{r})$, $C((\sigma_{i_{1}}^{\varepsilon_{1}}B\sigma_{i_{1}}^{-\varepsilon_{1}})^{r})$, $C((\sigma_{i_{2}}^{\varepsilon_{2}}\sigma_{i_{1}}^{\varepsilon_{1}}B\sigma_{i_{1}}^{-\varepsilon_{1}}\sigma_{i_{2}}^{-\varepsilon_{2}})^{r})$, $\dots$ are similar to each other when $B^{r}$ is pure by Lemma 2, we have the following corollary.
Corollary 2.
If $B^{\prime}\in\mathscr{B}_{n}$ is conjugate to $B\in\mathscr{B}_{n}$, then $C(B^{r})$ and $C((B^{\prime})^{r})$ are similar to each other, where $r$ is the order of the permutation of $B$.
@@
Take braid diagrams of $b$ and $b^{\prime}$ in the form of $B$ and $B^{\prime}=A^{-1}BA\in\mathscr{B}_{n}$. By Remark 1 and Corollary 2, the crossing matrices $C(b^{r})=C(B^{r})$ and $C((b^{\prime})^{r})=C((B^{\prime})^{r})$ are similar to each other. ∎
@@ 4 Definition of the purified determinant
In this section, we define the purified determinant.
@@
We note that $B^{|\rho|}$ is a pure braid diagram that is obtained by taking the product of $|\rho|$ braid diagrams $B$. The crossing number of $B^{|\rho|}$ may be large. The following proposition implies that we can also calculate the value of $P(B)$ by considering the crossing matrix of $B$ instead of $B^{|\rho|}$.
@@ 5 Proof of the main theorem and examples
In this section, we prove Theorem 1 and see examples.
Proof of Theorem 1. If $b$ and $b^{\prime}\in\mathfrak{B}_{n}$ are conjugate, then the crossing matrices $C(b^{r})$ and $C((b^{\prime})^{r})$ are similar to each other by Proposition 3, where $r$ is the order of the braid permutation of $b$ and $b^{\prime}$. Hence, their determinants are equal, namely, $P(b)=P(b^{\prime})$. ∎
@@
The rank, eigenvalues [^1] It is shown in [5] that the crossing matrix of a pure braid is symmetric. Hence, the matrix $C(b^{|\rho|})$ definitely has real eigenvalues for any braid $b$., eigenvectors, and characteristic polynomial of $C(b^{|\rho|})$ are also braid invariants that are unchanged under conjugation.
@@
The braid diagrams $B$ and $B^{\prime}\in\mathscr{B}_{3}$ in Figure 4 represent different braids since they have different crossing matrices. Moreover, they are not conjugate since they have the different values of purified determinants as P(B)=8P(B)=-8 and $P(B^{\prime})=8$, even the order of their permutations are the same and the linking number of the closures (see the next section) are the same.
@@
For a braid diagram $B$ and its permutation $\rho$, let $PP(B)=(|\rho|,P(B))$. We call $PP(B)$ the P pair of $B$.
@@
The P pair is a braid invariant. Thai is, we can define the P pair of a braid $b\in\mathfrak{B}_{n}$ as $PP(b)=PP(B)$ for any diagram $B$ of $b$. Moreover, $PP(b)$ is unchanged under conjugation.
@@
Let $B\in\mathscr{B}_{3}$ be a braid diagram represented by $\sigma_{1}^{-1}\sigma_{2}$. Although $B$ and $B^{3}$ have the same value of the purified determinant by definition, they have different P pairs as $PP(B)=(3,0)$ and $PP(B^{3})=(1,0)$.
@@ 6 Future work
For an $n$-braid diagram $B\in\mathscr{B}_{n}$, a stabilization is the deformation to obtain $B\sigma_{n}^{\varepsilon}\in\mathscr{B}_{n+1}$, where $\varepsilon\in\{+1,-1\}$. The opposite deformation is called a destabilization. Markov showed in [8] that for two braids $b$ and $b^{\prime}$, the closures $\widehat{b}$ and $\widehat{b^{\prime}}$ represent the same link type if and only if $b$ and $b^{\prime}$ are related by a finite number of conjugations, stabilizations and destabilizations. (See, for example, [4], [7].) We note that the purified determinant can be changed by stabilization. For example, we have $P(B)=0$ and $P(B^{\prime})=-3$ for the braid diagrams in Figure 5.
@@
We remark that the Q-determinant can be changed under conjugations. For example, we have $Q(\sigma_{2}^{-3}\sigma_{1})=-1$ and $Q(\sigma_{2}^{-2}\sigma_{1}\sigma_{2}^{-1})=1$. Employing the crossing matrix, constructing a new braid invariant that is unchanged under both conjugations and stabilizations would be our future goal. It would also be a link invariant which distinguishes oriented or unoriented links from a new aspect.

### v3

@@ Abstract
In this paper, we define a braid invariant, the purified determinant $P(b)$ of a braid $b$, considering the determinant of the crossing matrix of a pure braid derived from $b$, and show that $P(b_{1}b_{2})=P(b_{2}b_{1})$ for any pair of $n$-braids $b_{1}$ and $b_{2}$.
@@ 1 Introduction
In this paper, we introduce a new braid invariant, the purified determinant, $P(b)$, of a braid $b\in\mathfrak{B}_{n}$ considering the determinant of the crossing matrix of a pure braid derived from $b$, and show that the value of $P$ is unchanged by a conjugation.
@@ 2 Crossing matrix
In this section, we review the crossing matrix and its properties. For a braid diagram, we call $\sigma_{i}$ a positive crossing and $\sigma_{i}^{-1}$ a negative crossing. For a matrix $M$, we denote the $(i,j)$-component by $M(i,j)$. The crossing matrix $C(B)=M$ of an $n$-braid diagram $B$ was defined by Burillo, Gutierrez, Krstić and Nitecki in [5] (see also [6]) to be the $n\times n$ matrix such that $M(i,i)=0$ and $M(i,j)$ is the number of positive crossings minus negative crossings between the $i^{th}$ and $j^{th}$ strands where the $i^{th}$ strand is over the $j^{th}$ strand. (See Figure 2.)
@@
For a braid $b\in\mathfrak{B}_{n}$ or a braid diagram $B\in\mathscr{B}_{n}$, when a strand has the initial point at the $i^{th}$ position from the left and has the end point at the $j^{th}$ position from the left, we denote $\rho(i)=j$ and $\rho(1,2,\dots,n)=(j_{1},j_{2},\dots,j_{n})$ for $\rho(k)=j_{k}$. We call $\rho$ the braid permutation of $b$ or $B$. For example, the braid diagram $B$ in Figure 2 has the braid permutation $\rho$ as $\rho(1,2,3,4,5)=(2,1,4,5,3)$. We note that a braid and its diagram have the same braid permutation. By regarding $\rho$ as an element of the symmetric group, the identity $\rho^{0}=id$, the inverse $\rho^{-1}$, the exponentiation $\rho^{k}$ for $k\in\mathbb{N}$ are also defined. For a permutation $\rho$, the order, $|\rho|$, of $\rho$ is defined to be the number $k\in\mathbb{N}$ such that $\rho^{k}=id$ and $\rho^{l}\neq id$ for any $l<k$. We say that a braid or a braid diagram is pure when the permutation is identity. The following proposition was shown in [5].
Proposition 2 ([5]).
An $n\times n$ matrix $M$ is the crossing matrix of some pure $n$-braid if and only if $M$ is a zero-diagonal symmetric integer matrix [^1] It has been conjectured that an $n\times n$ integer matrix $M$ is the crossing matrix of some positive pure $n$-braid if and only if $M$ is a non-negative integer zero-diagonal symmetric matrix with a property “T0” ([5]), and addressed in [6], [9], [11]. .
We define an operation on a square matrix that reorders the row and column simultaneously according to a permutation $\rho$.
@@
As mentioned in [10], $M$ and $\rho(M)$ are similar, that is, there exists an invertible matrix $Q$ such that $\rho(M)=Q^{-1}MQ$, and therefore $\det(M)=\det(\rho(M))$. For two $n\times n$ matrices $M$ and $N$, we have $\rho(M+N)=\rho(M)+\rho(N)$ by definition. The following proposition was shown in [5] for the crossing matrix of a braid product.
@@ 3 Similarity of crossing matrices
Proof of Lemma 1. Let $M_{1}=C(B)$ and $M_{2}=C(B^{\prime})$, where $B^{\prime}=\sigma_{i}^{\varepsilon}B\sigma_{i}^{-\varepsilon}$. For $B^{\prime}$, the $i^{th}$ (respectively, $(i+1)^{th}$) strand corresponds to the $(i+1)^{th}$ (respectively, $i^{th}$) strand of $B$ because there is a crossing $\sigma_{i}^{\varepsilon}$ of $B^{\prime}$ between the $i^{th}$ and $(i+1)^{th}$ strands as the first crossing, which is above $B$. Also, the first crossing $\sigma_{i}^{\varepsilon}$ and the last crossing $\sigma_{i}^{-\varepsilon}$ of $B^{\prime}$ are both between the $i^{th}$ and $(i+1)^{th}$ strands because $B^{\prime}$ is pure, where the $(i+1)^{th}$ (respectively, $i^{th}$) strand is over when $\varepsilon=+1$ (respectively, $\varepsilon=-1$), and one of the two crossings is positive and the other is negative. (See Figure 3.) Hence, we have $M_{2}(i+1,i)=M_{1}(i,i+1)$ and $M_{2}(i,i+1)=M_{1}(i+1,i)$, and $M_{2}=\tau_{i,i+1}(M_{1})$, where $\tau_{i,i+1}$ is a permutation that swaps the $i^{th}$ and $(i+1)^{th}$ elements. This implies that $M_{1}$ and $M_{2}$ are similar to each other. $\square$
@@
Lemma 1 and Corollary 1 do not hold for non-pure braid diagrams. For example,
$C(\sigma_{1})=\begin{bmatrix}0&0&0\\ 1&0&0\\ 0&0&0\end{bmatrix}\text{ \ and }\ C(\sigma_{2}^{-1}\sigma_{1}\sigma_{2})=\begin{bmatrix}0&0&0\\ 1&0&-1\\ 1&0&0\end{bmatrix}$
are not similar. Observe that the crossing matrices of them to the second power
$C((\sigma_{1})^{2})=\begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{bmatrix}\text{ \ and }\ C((\sigma_{2}^{-1}\sigma_{1}\sigma_{2})^{2})=\begin{bmatrix}0&0&1\\ 0&0&0\\ 1&0&0\end{bmatrix}$
are similar to each other. We have the following lemma for non-pure braid diagrams.
@@
If $B^{\prime}\in\mathscr{B}_{n}$ is conjugate to $B\in\mathscr{B}_{n}$, then $C(B^{r})$ and $C((B^{\prime})^{r})$ are similar to each other, where $r$ is the order of the braid permutation of $B$.
@@
Let $B^{\prime}=A^{-1}BA$. Since $B^{r}$ is pure and $(B^{\prime})^{r}=(A^{-1}BA)(A^{-1}BA)\dots(A^{-1}BA)$ has the same crossing matrix to $A^{-1}B^{r}A$ by Proposition 1, $B^{r}$ and $(B^{\prime})^{r}$ have similar crossing matrices by Corollary 1. ∎
@@
Take braid diagrams of $b$ and $b^{\prime}$ in the form of $B$ and $B^{\prime}=A^{-1}BA\in\mathscr{B}_{n}$. By Proposition 1 and Lemma 2, the crossing matrices $C(b^{r})=C(B^{r})$ and $C((b^{\prime})^{r})=C((B^{\prime})^{r})$ are similar to each other. ∎
@@ 4 Definition of the purified determinant
In this section, we define the purified determinant and see examples.
@@
We note that $B^{|\rho|}$ is a pure braid diagram that is obtained by taking the product of $|\rho|$ braid diagrams $B$. The number of crossings of $B^{|\rho|}$ may be large. The following proposition implies that we can also calculate the value of $P(B)$ by considering the crossing matrix of $B$ instead of $B^{|\rho|}$.
@@
Example 3.
A fundamental braid, $\Delta_{n}$, is a positive $n$-braid that is obtained from a trivial braid by turning over the lower bar. We have
$P(\Delta_{n})=-(-1)^{n}(n-1).$
Proof.
The order of permutation of each $\Delta_{n}$ is 2, and each $\Delta_{n}$ has an $n$-braid diagram such that each pair of strands has exactly one positive crossing. Hecne, the matrix $M_{n}=C((\Delta_{n})^{2})$ is a matrix such that $M_{n}(i,i)=0$ and $M_{n}(i,j)=1$ for $i\neq j$. To calculate $\det(M_{n})$, subtract the $2^{nd}$ column from the $1^{st}$ column, and apply a Laplace expansion along the $1^{st}$ column. Then, break down the $1^{st}$ row of the $2^{nd}$ term, and we obtain a recurrence formula $\det(M_{n})=-2\det(M_{n-1})-\det(M_{n-2})$. For example, when $n=5$, we have
$\begin{vmatrix}0&1&1&1&1\\ 1&0&1&1&1\\ 1&1&0&1&1\\ 1&1&1&0&1\\ 1&1&1&1&0\end{vmatrix}=\begin{vmatrix}-1&1&1&1&1\\ 1&0&1&1&1\\ 0&1&0&1&1\\ 0&1&1&0&1\\ 0&1&1&1&0\end{vmatrix}=-\begin{vmatrix}0&1&1&1\\ 1&0&1&1\\ 1&1&0&1\\ 1&1&1&0\end{vmatrix}-\begin{vmatrix}1&1&1&1\\ 1&0&1&1\\ 1&1&0&1\\ 1&1&1&0\end{vmatrix}$
$=-\begin{vmatrix}0&1&1&1\\ 1&0&1&1\\ 1&1&0&1\\ 1&1&1&0\end{vmatrix}-\left(\begin{vmatrix}0&1&1&1\\ 1&0&1&1\\ 1&1&0&1\\ 1&1&1&0\end{vmatrix}+\begin{vmatrix}1&0&0&0\\ 1&0&1&1\\ 1&1&0&1\\ 1&1&1&0\end{vmatrix}\right)=-2\begin{vmatrix}0&1&1&1\\ 1&0&1&1\\ 1&1&0&1\\ 1&1&1&0\end{vmatrix}-\begin{vmatrix}0&1&1\\ 1&0&1\\ 1&1&0\end{vmatrix}$.
@@ 5 Proof of the main theorem and examples
In this section, we prove Theorem 1.
Proof of Theorem 1. If $b$ and $b^{\prime}\in\mathfrak{B}_{n}$ are conjugate, then the crossing matrices $C(b^{r})$ and $C((b^{\prime})^{r})$ are similar to each other by Proposition 4, where $r$ is the order of the braid permutation of $b$ and $b^{\prime}$. Hence, their determinants are equal, namely, $P(b)=P(b^{\prime})$. $\square$
@@
The rank, eigenvalues, and characteristic polynomial of $C(b^{|\rho|})$ are also braid invariants that are unchanged under conjugation.
We note that the matrix $C(b^{|\rho|})$ definitely has real eigenvalues for any braid $b$ since the crossing matrix of the pure braid $b^{|\rho|}$ is symmetric by Proposition 1.
@@
The braid diagrams $B$ and $B^{\prime}\in\mathscr{B}_{3}$ in Figure 4 represent different braids since they have different crossing matrices. Moreover, they are not conjugate since they have the different values of purified determinant [^2] Eigenvalues of $C(B^{2})$ are $1$, $\frac{-1\pm\sqrt{33}}{2}$ and that of $C((B^{\prime})^{2})$ are 1-1, $\frac{1\pm\sqrt{33}}{2}$. This also can conclude that $B$ and $B^{\prime}$ are not conjugate. as P(B)=8P(B)=-8 and $P(B^{\prime})=8$, even the order of their permutations are the same and the linking number of the closures (see Section 6) are the same.
@@
A braid diagram $B$ is said to be layered when the set of all the strands of $B$ can be divided into two sets $S_{1}$ and $S_{2}$ so that any strand in $S_{1}$ does not pass under a strand in $S_{2}$. We denote a braid diagram that is obtaind from $B$ by ignoring all the strands in $S_{2}$ (respectively, $S_{1}$) by $B_{1}$ (respectively, $B_{2}$) and denote as $B=B_{1}\oplus B_{2}$. For pure braid diagrams, we have the following proposition in the same way to Theorem 1 in [10].
Proposition 6.
When a layered braid diagram $B=B_{1}\oplus B_{2}$ is pure, we have $P(B)=P(B_{1})P(B_{2})$.
Proof.
Take a sequence $\mathbf{s}=(s^{1}_{1},s^{1}_{2},\dots,s^{1}_{k_{1}},s^{2}_{1},s^{2}_{2},\dots,s^{2}_{k_{2}})$ of the strands of $B$ so that $s^{1}_{1},s^{1}_{2},\dots,s^{1}_{k_{1}}\in S_{1}$ and $s^{2}_{1},s^{2}_{2},\dots,s^{2}_{k_{2}}\in S_{2}$. Reorder the row and column of $C(B)$ simultaneously according to the order of $\mathbf{s}$ to obtain a matrix $M$, which is similar to $C(B)$. Observe that $M$ is in the form of
$M=\begin{bmatrix}M_{1}&O\\ O&M_{2}\end{bmatrix}$
with the blocks $M_{1}=C(B_{1})$ and $M_{2}=C(B_{2})$. We note that $C(B)$ is symmetric by Proposition 1 and $M$ is also symmetric. The block on the lower left is $O$ because no strand in $S_{2}$ passes over a strand in $S_{1}$. The block on the upper right is also $O$ by symmetry.
Since $B$ is pure, we have $P(B)=\det(C(B))$, and $\det(C(B))=\det(M)=\det(M_{1})\det(M_{2})=P(B_{1})P(B_{2})$. We note that $B_{1}$ and $B_{2}$ are also pure, and therefore $P(B_{i})=\det(C(B_{i}))=\det(M_{i})$. ∎
@@
For a braid diagram $B$ and its permutation $\rho$, let $PP(B)=(|\rho|,P(B))$. We call $PP(B)$ the P-pair of $B$.
@@
The P-pair is a braid invariant. Thai is, we can define the P pair of a braid $b\in\mathfrak{B}_{n}$ as $PP(b)=PP(B)$ for any diagram $B$ of $b$. Moreover, $PP(b)$ is unchanged under conjugation.
@@
Let $B\in\mathscr{B}_{3}$ be a braid diagram represented by $\sigma_{1}^{-1}\sigma_{2}$. Although $B$ and $B^{3}$ have the same value of the purified determinant by definition, they have different P-pairs as $PP(B)=(3,0)$ and $PP(B^{3})=(1,0)$.
@@ 6 Future work
For an $n$-braid diagram $B\in\mathscr{B}_{n}$, a stabilization is the deformation to obtain $B\sigma_{n}^{\varepsilon}\in\mathscr{B}_{n+1}$, where $\varepsilon\in\{+1,-1\}$. The opposite deformation is called a destabilization. Markov showed in [8] that for two braids $b$ and $b^{\prime}$, the closures $\widehat{b}$ and $\widehat{b^{\prime}}$ represent the same link type if and only if $b$ and $b^{\prime}$ are related by a finite number of conjugations, stabilizations and destabilizations. (See, for example, [4], [7].) We note that the purified determinant can be changed by a stabilization. For example, we have $P(B)=0$ and $P(B^{\prime})=-3$ for the braid diagrams in Figure 5.
@@
We note that the Q-determinant can be changed under conjugations. For example, we have $Q(\sigma_{2}^{-3}\sigma_{1})=-1$ and $Q(\sigma_{2}^{-2}\sigma_{1}\sigma_{2}^{-1})=1$. Employing the crossing matrix, constructing a new braid invariant that is unchanged under both conjugations and stabilizations would be our future goal. It would also be a link invariant which distinguishes oriented or unoriented links from a new aspect.